CAM CHIP HIERARCHY

u1 cam
u11 data_path
u8 glue_decode
u1 glue_reg
u6 glue_reg
u7 glue_reg
u8 glue_reg
u9 glue_reg
u10 glue_reg
u22 glue_xin_mux
u23 glue_yin_mux
u24 glue_zin_mux
u2 glue_output_mux
u100 glue_pipe
u12 ssr_decode
u344 src_mxr
u345 src_mxr
u346 src_mxr
u347 src_mxr
u348 src_mxr
u80 src_delay
u10 evc_control
u1 cntr5_dn
u9 event_ctr
u274 cntr8_src
u275 cntr8_src
u278 cntr8_src
u192 par_ser
u193 par_ser
u194 par_ser
u11 dram_cnt_reg
u104 cntr8_ud_sr
u105 cntr8_ud_sr
u10 scan_gen
u1 site_gen
u1 cntr8_sr
u2 cntr8_sr
u3 cntr8_sr
u159 mxr24_1
u230 mxr24_1a
u3 kick_reg
u281 kick_carry

u2 offset_reg
u192 par_ser
u193 par_ser
u194 par_ser
u4 dim_reg
u280 mxr24_1b
u281 mxr24_1b
u282 mxr24_1b
u6 run_reg
u167 sync_dcd
u3 cntr6_dn
u180 sc_latch
u5 scan_reg
u200 bex35
u122 mux24_1
u33 sram_if
u23 lutmac_reg
u83 cntr_sr8
u84 cntr_sr8
u40 dram_if
u90 nibble_reg
u91 nibble_fifo
u85 addr_pipe
u99 scan
u3 scan1_p_mksc
u101 scan_dly
u271 est_dly
u100 dram_ctrl
u40 dram_ctrl_p_mksc
u9 cntr5_dn
u10 cntr8_dn
u13 busif
u15 opc_dcd
u16 opc_reg
u17 reg_len_dcd
u52 cntr7_dn
u200 reg_len_p_mksc
u18 reg_out_mux
u313 mux29_1
u20 hw_env_reg
u34 mpc_reg_a
u35 mpc_reg_b
u36 vwe_reg
u19 interrupt_reg

u328 ser_latch
u14 node
u34 newnodeio
rwfsm_p
node_drc
nodefsm_p
node_inv
node_sel